OpenStudy (anonymous):

I don't know how to type this in the correct form but its the fifth root of (5x^7)(y^2)/(27x^2)

OpenStudy (anonymous):

\[ \sqrt[5]{5^{7}y^{2}/27x^{2}}\]

OpenStudy (anonymous):

something like that

OpenStudy (anonymous):

between the 5 and the ^{7} is suppose be an x

OpenStudy (anonymous):

\[\sqrt[5]{\frac{5x^7y^2}{27x^2}}=\sqrt[5]{\frac{5x^5y^2}{27}}=x \sqrt[5]{\frac{5y^2}{27}}\]